OpenStudy (anonymous):

Which of the following statements have the same result? Explain each step in solving each one. I.f(1) when f(x) = 5x + 1 II.f-1(3) when f(x) = 2x plus 3, all over 5 III.3y - 7 = y + 5 @loser66

OpenStudy (anonymous):

@kc_kennylau @klorainewilliams

OpenStudy (anonymous):

@HazelLuv99

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

At least i. and iii. yield the same result, 6.

OpenStudy (anonymous):

what? @robtobey I have no idea how to do this..

OpenStudy (anonymous):

i. Evaluate f(x) at x=1, 5*1 + 1 = 6 iii. 3y - 7 = y + 5 Solve the above equation for y. y = 6